# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v11

@@ 5.2 Theorem of Ethical Safety Assurance
Definition 5 (Ethical Safety) Let $S(X,T)$ be the interpretation of a set of well-formed formulas in a formal system $\mathcal{L}$ over the domain $X\times T$. Let $Ec(\overrightarrow{v})$ denote an ethical constraint expressed as a well-formed formula in $\mathcal{L}$, where $\overrightarrow{v}=(v_{1},\ldots,v_{n})$ are its free variables. For each substitutable $(x,t)\in X\times T$, define
@@
$\varphi_{u}\coloneqq$ $\bigl((\forall\,\Phi_{u})(\forall\,\Phi_{q})\bigl({IsSubsetOf}^{2}(\Phi_{u},S_{ou}(ou,{T}_{u,h}))$ (5.4)
$\quad\land{IsSubsetOf}^{2}(\Phi_{q},S_{oq}(oq,{T}_{q,h}))$
$\quad\land{Processable}^{2}(\Phi_{u},\Phi_{q})\bigr)\bigr)$
$\rightarrow\bigl(\exists\,\varphi_{r}\bigl({IsFormulaOf}^{2}(\varphi_{r},S_{or}(or,{T}_{r,h}))$
$\quad\land{Eq}^{2}(\varphi_{r},f_{\mathrm{process}}^{2}(\Phi_{u},\Phi_{q}))\bigr)\bigr).$
@@
$\begin{split}S_{or}\left(or,{Tr}_{h}\right)=\left\{\varphi r\middle|\right.&\left.\varphi r={Processable}^{2}(\Phi u,\Phi q),\right.\\ &\left.\Phi u\in Sou\left(ou,{Tu}_{h}\right),\right.\\ &\left.\Phi q\in S_{oq}\left(oq,{Tq}_{h}\right)\right\}\end{split}$ (5.5)
@@
where the vertex set is precisely $S_{or}\left(or,{Tr}_{h}\right)$, and the hyperedge set $\mathcal{E}$ covers all well-formed formulas that directly or implicitly violate the ethical constraint $Ec$:
@@
By the maximum independent set theorem for finite hypergraphs, $H_{or}$ necessarily contains an independent set $G_{\max}$ of maximal cardinality. Suppose there exists another ethically safe subset $K$ with
@@
In this way, add to $Sou\left(ou,{Tu}_{h}\right)$ an ethical safeguard formula:
$\begin{split}\varphi u=&(\forall\varphi r)({IsFormulaOf}^{2}\left(\varphi r,S_{or}\left(or,{Tr}_{h}\right)\right))\rightarrow\exists\varphi s\left({IsFormulaOf}^{2}\left(\varphi s,S_{os}\left(os,{Ts}_{h}\right)\right)\right.\\ &\left.\land\left({IsFormulaOf}^{2}\left(\varphi r,S_{or}^{Ec}\left(or,{Tr}_{h}\right)\right){\rightarrow Eq}^{2}(\varphi s,\varphi r)\right)\right.\\ &\left.\land\left({IsFormulaOf}^{2}\left(\varphi r,S_{or}^{\sim Ec}\left(or,{Tr}_{h}\right)\right){\rightarrow Eq}^{2}\left(\varphi s,{IsSafeback}^{1}(NEc)\right)\right)\right)\end{split}$ (5.9)
Since the newly added $\varphi u$ pertains to $S_{os}\left(os,{Ts}_{h}\right)$ and is entirely independent of the existing formulas in $Sou\left(ou,{Tu}_{h}\right)$, it does not affect the satisfaction of Postulate 5 by $Sou\left(ou,{Tu}_{h}\right)$. Moreover,
@@
By taking $S_{os}\left(os,{Ts}_{h}\right)$ as the ontological state and mapping it via recoverable information $Is$ to the carrier state $S_{M}(M,{Ts}_{m})$, $S_{M}(M,{Ts}_{m})$ is necessarily also ethically safe with respect to $Ec$. This satisfies the condition in Definition 5 for model M to be ethically safe with respect to $Ec$. The theorem is thus proved.
Theorem 2 not only establishes that it is possible to construct an ethical safety assurance mechanism for models, but also provides a constructive algorithm—based on hypergraphs—for obtaining the maximum ethically safe output. However, computing the maximum independent set in a hypergraph is generally an NP-hard problem. In practical applications, the algorithm can be simplified by leveraging two functions: one for ranking the conflict intensity among formulas in $S_{or}\left(or,{Tr}_{h}\right)$, and another for ethical safety checking, thus yielding an approximately maximal safe output. This is a practical issue that warrants further investigation and will not be elaborated upon here. Meanwhile, the inheritance of ethical safety in the learning process can be achieved by incorporating dedicated ethical constraint regularization terms into the model algorithm. This involves more specific model structures and algorithmic design, which are beyond the scope of this paper.
In summary, based on the MLT-MF framework, Definition 10 and Theorem 5, for the first time, construct a formal theoretical framework for ethical safety. This approach, unlike those based on heuristic rules or statistical indicators, fills the gap in mathematical foundations and verifiability in traditional research, and provides concrete methodological guidance for the automated embedding and verification of ethical constraints.
@@ 5.3 Theorem on TVD Upper Bounds
Applying the MLT-MF framework allows us to estimate upper bounds on the generalization error measured by Total Variation Distance (TVD) between model distributions, facilitating the analysis and prediction of model generalization performance across a variety of tasks.
Theorem 3 (TVD Upper Bound for Conservative Prediction Strategies in Noise-Free Inputs and Uncovered Training Regions) Let the information describing the model $M$ after learning and optimization be $Iu=\left\langle ou,{Tu}_{h},S_{ou},M,{Tu}_{m},S_{M}\right\rangle$, and the input information be $Iq=\left\langle oq,{Tq}_{h},S_{oq},cq,{Tq}_{m},S_{cq}\right\rangle$. Assume both $S_{ou}$ and $S_{oq}$ are distributed according to probability distributions over the same measurable space. Let $\left|S_{oq}\right|$ and $\left|S_{ou}\cap S_{oq}\right|$ denote the cardinalities of the finite sets $S_{oq}$ and $S_{ou}\cap S_{oq}$, respectively. The probability distribution of the model is $p_{M}(x)$, where $p_{M}(x)>0$ for $x\in S_{ou}$ ; the distribution over $S_{oq}$ is $p_{q}(x)$, with $p_{q}(x)>0$ for $x\in S_{oq}$. The total probability mass of $p_{q}(x)$ over $S_{ou}\cap S_{oq}$ is denoted $P_{Mq}$. When the model $M$ is sufficiently trained on $S_{ou}$, and adopts a conservative uniform prediction for unknown states, the total variation distance between $p_{M}(x)$ and $p_{q}(x)$ satisfies:
$TVD\left(p_{q},p_{M}\right)$ $=\begin{cases}\approx 0,&\text{if }S_{oq}\subset S_{ou}\\[5.0pt] \leq\sqrt{\frac{1}{2}((1-P_{Mq})\log{((\left|S_{oq}\right|-\left|S_{ou}\cap S_{oq}\right|)}/(1-P_{Mq}))-H(p_{q}))},&\text{if }S_{oq}\neq S_{ou}\cap S_{oq}\end{cases}$ (5.11)
@@
Using KL-divergence to measure the difference between the true distribution and the hypothesized model distribution, we have (Goodfellow et al., 2016):
@@
Since both $p_{M}(x)$ and $p_{q}(x)$ are probability distributions defined on the same measurable space, we can use KL-divergence to estimate an upper bound on their total variation distance. By Pinsker’s inequality (Pinsker, 1964), we have:
@@
Since we have already assumed that the carrier states $S_{M}(M,{Tu}_{m})$ and $S_{cq}(cq,{Tq}_{m})$ of $Iu=\left\langle ou,{Tu}_{h},S_{ou},M,{Tu}_{m},S_{M}\right\rangle$ and $Iq=\left\langle oq,{Tq}_{h},S_{oq},cq,{Tq}_{m},S_{cq}\right\rangle$ are both represent the physical realizations of $S_{ou}$ and $S_{oq}$ in model $M$ and its interface, respectively. It follows that equation (5.11) characterizes the upper bound of the model’s TVD. This completes the proof.
Distinct from traditional generalization theories such as VC dimension and Rademacher complexity, Theorem 3, based on the MLT-MF framework, utilizes $P_{Mq}$ and $\left|S_{oq}\right|-\left|S_{ou}\cap S_{oq}\right|$—which reflect the overlap between test data and the model’s original information—as well as the entropy $H\left(p_{q}\right)$ of test data in the non-overlapping region, as evaluation metrics to provide an explicit quantitative estimation of the TVD. Since TVD reflects the degree of distributional alignment of the model, independent of specific models and tasks, it is generally applicable for analyzing training performance. Moreover, for classification tasks with 0-1 loss, the upper bound of the generalization error is equivalent to the TVD (Devroye et al., 2013), while for regression tasks, the upper bound of the generalization error is given by the product of the TVD and the Lipschitz loss (Devroye et al., 2013). Therefore, TVD serves as an indicator for estimating a model’s generalization error. In this way, Theorem 3 offers a novel perspective and practical guidance for model design and training data acquisition.
@@
Thus, the upper bound of the TVD for model $M$ with respect to noisy input information is given by $TVD\left(\widetilde{p_{q}},p_{M}\right)\leq\sqrt{\frac{\left[E_{fit}+P_{uncov}D_{KL}\left(r\parallel p_{0}\right)\right]}{2}}$ (5.17)
@@
$TVD\left(p_{q},\widetilde{p_{q}}\right)=(\sum_{x\in X}\left|p_{q}(x)-\widetilde{p_{q}}(x)\right|)/2$
$(\sum_{x\in S_{oq}^{-}}\left|p_{q}(x)-\widetilde{p_{q}}(x)\right|+\sum_{x\in S_{oq}^{+}}\left|p_{q}(x)-\widetilde{p_{q}}(x)\right|+\sum_{x\in X\backslash(S_{oq}^{-}\cup S_{oq}^{+})}\left|p_{q}(x)-\widetilde{p_{q}}(x)\right|)/2$
$(\sum_{x\in S_{oq}^{-}}\left|p_{q}(x)\right|+\sum_{x\in S_{oq}^{+}}\left|p^{+}(x)/(1-P_{l}+P_{a})\right|+\sum_{x\in X\backslash(S_{oq}^{-}\cup S_{oq}^{+})}\left|p_{q}(x)-p_{q}(x)/(1-P_{l}+P_{a})\right|)/2$
$(P_{l}+\frac{P_{a}}{1-P_{l}+P_{a}}+\frac{\left|P_{l}-P_{a}\right|\sum_{x\in X\backslash(S_{oq}^{-}\cup S_{oq}^{+})}{p_{q}(x)}}{1-P_{l}+P_{a}})/2=(P_{l}+\frac{P_{a}}{1-P_{l}+P_{a}}+\frac{\left|P_{l}-P_{a}\right|(1-P_{l})}{1-P_{l}+P_{a}})/2$
$\left\{\begin{array}[]{r}P_{l},\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ if\ P_{l}\geq P_{a},\\ \frac{P_{a}}{1-P_{l}+P_{a}},\ \ \ \ \ \ \ \ \ \ \ \ \ \ if\ P_{l}<P_{a}.\end{array}\right$.\ (5.22)
@@
$TVD\left(p_{q},p_{M}\right)\leq TVD\left(p_{q},\widetilde{p_{q}}\right)+TVD\left(\widetilde{p_{q}},p_{M}\right)$
$\leq\left\{\begin{array}[]{r}P_{l}+\sqrt{\frac{\left[E_{fit}+P_{uncov}D_{KL}\left(r\parallel p_{0}\right)\right]}{2}},\ \ \ \ \ \ \ \ if\ P_{l}\geq P_{a},\\ \frac{P_{a}}{1-P_{l}+P_{a}}+\sqrt{\frac{\left[E_{fit}+P_{uncov}D_{KL}\left(r\parallel p_{0}\right)\right]}{2}},\ if\ P_{l}<P_{a}.\end{array}\right$.\ (5.23)
@@ 6.1 Experiments on the TVD upper bound under noiseless input and conservative prediction strategy
(c) For all test samples from the non-overlapping region, a conservative uniform prediction strategy was used. For almost all samples (except a few outliers), their RMSE is less than $TVD\approx\sqrt{((1-0.4951)\ln{((10000-4951)}/(1-0.4951))-2.9741)/2}\approx 0.9155$. This indicates that even in regression tasks, although TVD is not strictly equivalent to generalization error, it can still effectively capture the distributional characteristics of generalization error.
@@
In summary, this regression experiment using a simple feedforward neural network model demonstrates and validates the accuracy of the model TVD upper bound estimation based on the MLT-MF framework, and its potential for estimating generalization error. Since the theoretical proof of Theorem 3 is independent of specific models and tasks, it promises to be extensible to more complex scenarios.
Figure 3: Experimental Evaluation and Results of the TVD Upper Bound under Noiseless Input and Conservative Prediction Strategy. (a) The blue rectangular region represents the training dataset comprising 10,000 samples, while the red circular region represents the test dataset, also with 10,000 samples. Among these, 4,951 samples overlap with the training set. (b) For each of the training steps (10, 50, 100, 200, 500, 1000, 4000), the variation in root mean square error (RMSE) is shown for 10 randomly selected test samples from the training dataset. (c) In the non-overlapping region between the test and training datasets, the RMSEs for all 5,049 test samples are compared between the conservative uniform prediction strategy and the TVD upper bound. (d) For 10 randomly selected test samples from the non-overlapping region, the changes in RMSE are shown for training steps (10, 30, 300, 600, 1000, 4000, 7000), along with comparisons to both the TVD upper bound and the RMSE from the conservative uniform prediction.

### v12

@@ 5.2 Theorem of Ethical Safety Assurance
Definition 5 (Ethical Safety) Let $S(X,T)$ be the interpretation of a set of WFFs in a formal system $\mathcal{L}$ over the domain $X\times T$. Let $Ec(\overrightarrow{v})$ denote an ethical constraint expressed as a WFF in $\mathcal{L}$, where $\overrightarrow{v}=(v_{1},\ldots,v_{n})$ are its free variables. For each substitutable $(x,t)\in X\times T$, define
@@
$\varphi u=$ $\bigl((\forall\,\Phi u)(\forall\,\Phi q)\bigl({IsSubsetOf}^{2}(\Phi u,S_{ou}(ou,{Tu}_{h}))$ (5.4)
$\quad\land{IsSubsetOf}^{2}(\Phi q,S_{oq}(oq,{Tq}_{h}))$
$\quad\land{Processable}^{2}(\Phi u,\Phi q)\bigr)\bigr)$
$\rightarrow\bigl(\exists\,\varphi_{r}\bigl({IsFormulaOf}^{2}(\varphi r,S_{or}(or,{Tr}_{h}))$
$\quad\land{Eq}^{2}(\varphi r,f_{\mathrm{process}}^{2}(\Phi u,\Phi q))\bigr)\bigr).$
@@
$\begin{split}S_{or}\left(or,{Tr}_{h}\right)=\left\{\varphi r\middle|\right.&\left.\varphi r=f_{\mathrm{process}}^{2}(\Phi u,\Phi q),\right.\\ &\left.\Phi u\in Sou\left(ou,{Tu}_{h}\right),\right.\\ &\left.\Phi q\in S_{oq}\left(oq,{Tq}_{h}\right)\right\}\end{split}$ (5.5)
@@
where the vertex set is precisely $S_{or}\left(or,{Tr}_{h}\right)$, and the hyperedge set $\mathcal{E}$ covers all WFFs that directly or implicitly violate the ethical constraint $Ec$:
@@
By the maximum independent set theorem for finite hypergraphs[36], $H_{or}$ necessarily contains an independent set $G_{\max}$ of maximal cardinality. Suppose there exists another ethically safe subset $K$ with
@@
In this way, add to $S_{ou}\left(ou,{Tu}_{h}\right)$ an ethical safeguard formula:
$\begin{split}\varphi u=&(\forall\varphi r)\left(\mathit{IsFormulaOf}^{2}\left(\varphi r,S_{or}\left(or,Tr_{h}\right)\right)\right)\rightarrow\\ &\exists\varphi s\left(\mathit{IsFormulaOf}^{2}\left(\varphi s,S_{os}\left(os,Ts_{h}\right)\right)\land\right.\\ &\left(\mathit{IsFormulaOf}^{2}\left(\varphi r,S_{or}^{Ec}\left(or,Tr_{h}\right)\right)\rightarrow\mathit{Eq}^{2}(\varphi s,\varphi r)\right)\land\\ &\left.\left(\mathit{IsFormulaOf}^{2}\left(\varphi r,S_{or}^{\sim Ec}\left(or,Tr_{h}\right)\right)\rightarrow\mathit{Eq}^{2}\left(\varphi s,\mathit{IsSafeback}^{1}(NEc)\right)\right)\right)\end{split}$ (5.9)
Since the newly added $\varphi u$ pertains to $S_{os}\left(os,{Ts}_{h}\right)$ and is entirely independent of the existing formulas in $S_{ou}\left(ou,{Tu}_{h}\right)$, it does not affect the satisfaction of Postulate A.1 by $S_{ou}\left(ou,{Tu}_{h}\right)$. Moreover,
@@
By taking $S_{os}\left(os,{Ts}_{h}\right)$ as the ontological state and mapping it via noise-free information $Is$ to the carrier state $S_{M}(M,{Ts}_{m})$, $S_{M}(M,{Ts}_{m})$ is necessarily also ethically safe with respect to $Ec$. This satisfies the condition in Definition 5 for model $M$ to be ethically safe with respect to $Ec$. The theorem is thus proved.
Theorem 2 not only establishes that it is possible to construct an ethical safety assurance mechanism for models, but also provides a constructive algorithm—based on hypergraphs—for obtaining the SafeMax. However, computing the maximum independent set in a hypergraph is generally an NP-hard problem. In practical applications, the algorithm can be simplified by leveraging two functions: one for ranking the conflict intensity among formulas in $S_{or}\left(or,{Tr}_{h}\right)$, and another for ethical safety checking, thus yielding an approximately SafeMax. This is a practical issue that warrants further investigation and will not be elaborated upon here. Meanwhile, the inheritance of ethical safety in the learning process can be achieved by incorporating dedicated ethical constraint regularization terms into the model algorithm. This involves more specific model structures and algorithmic design, which are beyond the scope of this paper.
@@ 5.3 Theorem on TVD Upper Bounds
Applying the MLT-MF framework allows us to estimate upper bounds on the generalization error measured by TVD between model distributions, facilitating the analysis and prediction of model generalization performance across a variety of tasks.
Theorem 3 (TVD Upper Bound for Conservative Prediction Strategies in Noise-Free Inputs and Uncovered Training Regions) Let the information describing the model $M$ after learning and optimization be $Iu:S_{ou}(ou,Tu_{h})\rightrightarrows S_{M}(M,Tu_{m})$, and the input information be $Iq:S_{oq}(oq,Tq_{h})\rightrightarrows S_{cq}(cq,Tq_{m})$. Assume both $S_{ou}$ and $S_{oq}$ are distributed according to probability distributions over the same measurable space. Let $|S_{oq}|$ and $|S_{ou}\cap S_{oq}|$ denote the cardinalities of the finite sets $S_{oq}$ and $S_{ou}\cap S_{oq}$, respectively. The probability distribution of the model is $p_{M}(x)$, where $p_{M}(x)>0$ for $x\in S_{ou}$; the distribution over $S_{oq}$ is $p_{q}(x)$, with $p_{q}(x)>0$ for $x\in S_{oq}$. The total probability mass of $p_{q}(x)$ over $S_{ou}\cap S_{oq}$ is denoted $P_{Mq}$. When the model $M$ is sufficiently trained on $S_{ou}$, and adopts a conservative uniform prediction for unknown states, the total variation distance (TVD) between $p_{M}(x)$ and $p_{q}(x)$ satisfies:
$\mathrm{TVD}(p_{q},p_{M})=\begin{cases}\approx 0,&\text{if }S_{oq}\subset S_{ou}\\[10.00002pt] \leq\sqrt{\dfrac{1}{2}D_{\text{KL}}},&\text{if }S_{oq}\neq S_{ou}\cap S_{oq}\end{cases}$ (5.11)
where $D_{\text{KL}}=(1\!-\!P_{\!Mq})\log\!\left(\!\dfrac{|S_{oq}|\!-\!|S_{ou}\cap S_{oq}|}{1\!-\!P_{\!Mq}}\!\right)\!-\!H(p_{q})$.
@@
Using KL-divergence to measure the difference between the true distribution and the hypothesized model distribution, we have [29]:
@@
Since both $p_{M}(x)$ and $p_{q}(x)$ are probability distributions defined on the same measurable space, we can use KL-divergence to estimate an upper bound on their TVD. By Pinsker’s inequality [30], we have:
@@
Since we have already assumed that the carrier states $S_{M}(M,Tu_{m})$ and $S_{cq}(cq,Tq_{m})$ of $Iu:S_{ou}(ou,Tu_{h})\rightrightarrows S_{M}(M,Tu_{m})$ and $Iq:S_{oq}(oq,Tq_{h})\rightrightarrows S_{cq}(cq,Tq_{m})$ are both represent the physical realizations of $S_{ou}$ and $S_{oq}$ in model $M$ and its interface, respectively. It follows that equation (5.11) characterizes the upper bound of the model’s TVD. This completes the proof.
Distinct from traditional generalization theories such as VC dimension and Rademacher complexity, Theorem 3, based on the MLT-MF framework, utilizes $P_{Mq}$ and $|S_{oq}|-|S_{ou}\cap S_{oq}|$—which reflect the overlap between test data and the model’s original information—as well as the entropy $H(p_{q})$ of test data in the non-overlapping region, as evaluation metrics to provide an explicit quantitative estimation of the TVD. Since TVD reflects the degree of distributional alignment of the model, independent of specific models and tasks, it is generally applicable for analyzing training performance. Moreover, for classification tasks with 0-1 loss, the upper bound of the generalization error is equivalent to the TVD [31], while for regression tasks, the upper bound of the generalization error is given by the product of the TVD and the Lipschitz loss [32]. Therefore, TVD serves as an indicator for estimating a model’s generalization error. In this way, Theorem 3 offers a novel perspective and practical guidance for model design and training data acquisition.
@@
Thus, the upper bound of the TVD for model $M$ with respect to noisy input information is given by
$TVD\left(\widetilde{p_{q}},p_{M}\right)\leq\sqrt{\frac{\left[E_{fit}+P_{uncov}D_{KL}\left(r\parallel p_{0}\right)\right]}{2}}$ (5.17)
@@
[…]
$=\frac{1}{2}\bigg(\sum_{x\in S_{oq}^{-}}\left|p_{q}(x)-\widetilde{p_{q}}(x)\right|+\sum_{x\in S_{oq}^{+}}\left|p_{q}(x)-\widetilde{p_{q}}(x)\right|$
$\quad+\sum_{x\in X\setminus(S_{oq}^{-}\cup S_{oq}^{+})}\left|p_{q}(x)-\widetilde{p_{q}}(x)\right|\bigg)$
$=\frac{1}{2}\bigg(\sum_{x\in S_{oq}^{-}}\left|p_{q}(x)\right|+\sum_{x\in S_{oq}^{+}}\left|\frac{p^{+}(x)}{1-P_{l}+P_{a}}\right|$
$\quad+\sum_{x\in X\setminus(S_{oq}^{-}\cup S_{oq}^{+})}\left|p_{q}(x)-\frac{p_{q}(x)}{1-P_{l}+P_{a}}\right|\bigg)$
$=\frac{1}{2}\left(P_{l}+\frac{P_{a}}{1-P_{l}+P_{a}}+\frac{|P_{l}-P_{a}|\sum_{x\in X\setminus(S_{oq}^{-}\cup S_{oq}^{+})}p_{q}(x)}{1-P_{l}+P_{a}}\right)$
$=\frac{1}{2}\left(P_{l}+\frac{P_{a}}{1-P_{l}+P_{a}}+\frac{|P_{l}-P_{a}|(1-P_{l})}{1-P_{l}+P_{a}}\right)$
$=\begin{cases}P_{l},&\text{if }P_{l}\geq P_{a},\\[2.0pt] \dfrac{P_{a}}{1-P_{l}+P_{a}},&\text{if }P_{l}<P_{a}.\end{cases}$ (5.24)
@@
$\text{TVD}(p_{q},p_{M})$ $\leq\text{TVD}(p_{q},\widetilde{p_{q}})+\text{TVD}(\widetilde{p_{q}},p_{M})$ (5.25)
[…]
@@ 6.1 Experiments on the TVD upper bound under noiseless input and conservative prediction strategy
(c) For all test samples from the non-overlapping region, a conservative uniform prediction strategy was used. For almost all samples (except a few outliers), their RMSE is less than $\mathrm{TVD}\approx\sqrt{((1-0.4951)\ln{((10000-4951)}/(1-0.4951))-2.9741)/2}\approx 0.9155$. This indicates that even in regression tasks, although TVD is not strictly equivalent to generalization error, it can still effectively capture the distributional characteristics of generalization error.
@@
Figure 3: Experimental Evaluation and Results of the TVD Upper Bound under Noiseless Input and Conservative Prediction Strategy.